OpenStudy (anonymous):

Solve the equation for x: 7x – 3y = 1

OpenStudy (anonymous):

ADD 3y divide by 7 7x – 3y = 1 7x=3y+1 x=(3/7)x+(1/7)

OpenStudy (anonymous):

Like this @zordoloom 

OpenStudy (anonymous):

can you draw? @zordoloom

OpenStudy (anonymous):